# Large Scale Knowledge Washing

### v2

@@ 6.2.1 Small-Scale Knowledge Washing
We first test the performances of our method on the small-scale knowledge-washing tasks, i.e., forgetting the knowledge in zsRE and CounterFactual. We report the results in Table 1. As shown in the table, our method can achieve the best performance concerning the cleanness of knowledge washing (measured by Accuracy and QA-F1-Score) while maintaining performance levels comparable to the original model on reasoning tasks. As the dataset scale is not large, it is shown in the table that WOH and SeUL, two fine-tuning-based methods achieve some certain extent of knowledge washing and also successfully maintain the model’s original ability, although there is already sign of performance degradation as shown in the dataset CounterFactual (See the performances of SeUL on GPT2-XL). Meanwhile, the method FT-UL could not achieve reasonable results as the reverse training objective is overly fragile to the training without more complicated regularization. We also show the generated examples for visualization in Appendix C.2.3.
@@ 6.2.2 Large Scale Knowledge Washing
Original 1.0000 0.3734 0.5121 10.63 1.0000 0.2553 0.6831 4.10
FT 0.0446 0.0256 0.1475 250.73 0.0159 0.0115 0.4349 13.00
MEMIT 0.2972 0.2342 0.4906 11.44 0.2536 0.0753 0.6817 4.32
ME-FT 0.0000 0.0000 0.0000 $>10^{10}$ 0.0000 0.0000 0.0000 $>10^{10}$
FT-UL 0.0000 0.0000 0.0000 $>10^{10}$ 0.0000 0.0000 0.0000 $>10^{10}$
WOH 0.4672 0.2227 0.1473 254.08 0.0009 0.0000 0.0171 $>10^{10}$
SeUL 0.0000 0.0000 0.0000 $>10^{10}$ 0.0004 0.0000 0.0002 $>10^{10}$
LaW 0.1926 0.1735 0.4657 13.27 0.1385 0.0846 0.6567 4.76
@@ 6.3 Ablation Study
Ablation Study on Initialization of $\hat{\Delta}$
We compare the performance of LaW on the dataset zsRE and CounterFactual with model GPT2-XL between using random initialization and using the initialization from MEMIT. For random initialization, we sample a matrix matching the dimension of $W_{0}$ (in Eq.(2)), filled with independent Gaussian random variables scaled by a factor of $0.001$: $\Delta_{0}=0.001\cdot\mathcal{N}(0,I)$. The results are reported in Table 3 (full table in Appendix C.2.2). When initializing from Gaussian distribution, we do not have reference $\beta_{0}$ as in the initialization from MEMIT, so we choose the constant $\beta=0.2$. Similarly, we also set $\beta=0.2$ when using MEMIT initialization. The table shows the MEMIT initialization can boost the performance drastically. The reason might be the optimization easily achieves local minimum when using random optimization.
Ablation Study of Choices of $\beta$
As shown in Eq.(10), the hyper-parameter $\beta$ can control the tradeoff between washing the knowledge in $K_{w}$ and maintaining the original knowledge in $K$ (which may also be related to the model’s reasoning ability, as we find that when this term is large the model’s reasoning ability may degrade drastically). To study the effects of different $\beta$, we choose the setting of dataset zsRE and CounterFactual with the model GPT2-XL to study the effects of different $\beta$. The results are reported in Table 4 (full table in Appendix C.2.2). From the table, we can see that as $\beta$ increases, the knowledge is washed more thoroughly and the reasoning abilities are also dropping, showing the tradeoff between knowledge washing and maintaining reasoning abilities. We can also find that setting $\beta$ according to $\beta_{0}$ can achieve better performances (see the performance comparison between $\beta=1.2\beta_{0}$ and $\beta=0.2$ on the dataset CounterFactual), which demonstrates the necessity of setting different $\beta$ for different layers.
Ablation Study on Successive Elimination of Knowledge Set
In our practical considerations, before modifying every layer, we find the facts in the knowledge set that the model can answer correctly and perform the knowledge washing on the selected knowledge set. To study the effects of this technique (denoted as SE), we conduct experiments with and without SE and report the results in Table 5 (full table in Appendix C.2.2). The results show that the algorithm can achieve a much cleaner washing with SE enabled, at the expense of slightly affecting the reasoning abilities.
@@
GPT2-XL 1.0000 0.3734 0.5121 10.63 1.0000 0.3734 0.5121 10.63
$\beta=1.05\beta_{0}$ 0.0074 0.0060 0.5127 10.74 0.1266 0.1070 0.4743 12.49
$\beta=1.1\beta_{0}$ 0.0050 0.0039 0.5108 10.86 0.1091 0.0905 0.4741 12.73
$\beta=1.2\beta_{0}$ 0.0008 0.0010 0.5073 11.06 0.0965 0.0853 0.4553 14.10
$\beta=1.5\beta_{0}$ 0.0000 0.0003 0.4945 12.02 0.0655 0.0587 0.4314 16.31
$\beta=0.1$ 0.0198 0.0166 0.5096 10.68 0.4318 0.2401 0.5063 10.82
$\beta=0.2$ 0.0008 0.0008 0.4628 14.34 0.1258 0.1102 0.4708 13.04
$\beta=0.5$ 0.0000 0.0000 0.2806 46.81 0.0242 0.0220 0.3169 36.23
Table 5: Ablation study with Successive Elimination technique enabled or disabled.
zsRE CounterFactual
Knowledge Reasoning Knowledge Reasoning
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 1.0000 0.3734 0.5121 10.63
LaW 0.0058 0.0043 0.5114 10.81 0.1091 0.0905 0.4741 12.73
w/o SE 0.0322 0.0254 0.5160 10.52 0.1345 0.1137 0.4905 12.20
@@ Appendix A Mathematical Details of Preliminary
As demonstrated by MEMIT [Meng et al., 2023], the objective is to adjust factual associations stored within the MLP layers of transformer-based, decoder-only large language models. The conditional distribution of the next token $x_{t}$, given by language model $G$, relies on the sequence of previous tokens:
@@
In MEMIT [Meng et al., 2023], $W_{out}^{l}$, denoted as $W_{0}$, can act as key-value memories, associating input keys $k_{i}\triangleq k_{i}^{l}$ with corresponding values $v_{i}\triangleq v_{i}^{l}$ [Geva et al., 2020]. If $W_{out}^{l}$ is dimensionally defined as $d_{1}\times d_{2}$ and stores $n$ memories, with $u$ new edits, then to modify the MLP layer $W_{out}^{l}$ (i.e., the matrix $W_{0}$), the following delta matrix $\Delta$ is solved:
@@
This process enables the editing of an MLP layer within the transformer $G$ to incorporate new relational data, following the solution of the equation for each $K_{e}$ and $V_{e}$ from the editing requests. In the MEMIT approach [Meng et al., 2023], $KK^{T}$ is pre-estimated and represented as $\lambda C_{0}$, where $C_{0}$ is the average covariance matrix of $K$ and $\lambda$ is a hyper-parameter typically on the order of 10,000.
When performing extensive model editing, modifying only one layer may lead to robustness issues, while a more stable model can be achieved by minimizing the magnitudes of parameter changes [Zhu et al., 2020]. Consequently, MEMIT proposes modifying multiple layers to distribute the editing impact more broadly [Meng et al., 2023]. This method involves spreading the residual $R=V_{e}-W_{0}K_{e}$ across several layers. Let $L$ represent the index of the deepest layer requiring modification such that the output of this layer transitions from $W_{0}K_{e}$ to $V_{e}$. Define $\mathcal{R}$ as the set of layer indices $\{L-|\mathcal{R}|+1,\ldots,L\}$ that require edits. For each layer $l$ within $\mathcal{R}$, the necessary adjustments to the weights $W^{l}$ are given by:
@@ Appendix B Implementation Details
For the baselines, we train GPT-J-6B with LoRA [Hu et al., 2021]. we put the configurations as below:
@@
    MEMIT. This method has a hyperparameter $\lambda$ when estimating $KK^{T}=\lambda C$ where $C$ is the average variable calculated on a large dataset (see the details in Meng et al. [2023]). The configurations of $\lambda$ in different settings are shown in Table 6. We found that with these configurations the model can achieve good knowledge-washing accuracy while mostly maintaining the model’s reasoning ability (minimal performance degradation on reasoning tasks.)
@@
    WOH. We first train the reinforced model on the sentences formed from the triplets $\mathcal{E}_{w}$ with the learning rate set as 1e-6 for 1 epoch, then we adopt the objective Eq.(1) from the paper Eldan and Russinovich [2023] to update the target model. During the second stage of training, we set the learning rate as 5e-5 and train the model for 1 epoch.
@@ Appendix C Additional Experiments
C.1 Descriptions of the Reasoning Datasets
We conduct the reasoning experiments on three datasets: Lambda_openai [Radford et al., 2019, Paperno et al., 2016], HellaSwag [Zellers et al., 2019], and Arc_Easy [Clark et al., 2018]. The descriptions of these three datasets are as follows:
@@ C.1 Descriptions of the Reasoning Datasets
    Lambda_openai [Paperno et al., 2016]: The LAMBADA dataset tests computational text understanding via a word prediction task. It features narrative texts where models must use broad context to predict the final word, rather than just the last sentence. The dataset includes an original test split and translations in German, Spanish, French, and Italian.
@@
    HellaSwag [Zellers et al., 2019]: The HellaSwag dataset is a benchmark designed for evaluating commonsense natural language inference (NLI) capabilities. It challenges models to complete sentences in a way that aligns with human common sense. The dataset prompts computational models to predict plausible sentence endings, testing their understanding of everyday scenarios and contexts.
@@
    ARC_Easy [Clark et al., 2018]: The ARC_Easy dataset is a subset of the ARC dataset, featuring grade-school level multiple-choice science questions that are less challenging compared to the full set. It includes questions that were correctly answered by standard algorithms.
C.2 Additional Experimental Results
C.2.1 Overall Performance Comparison
The overall performance comparisons with the performances on two other reasoning benchmarks on zsRE, CounterFactual and Wiki-Latest are shown in Table 10, Table 11 and Table 12, respectively.
C.2.2 Ablation Study
Ablation Study on Initialization of $\hat{\Delta}$
Ablation Study of Choices of $\beta$
We put the full results of different choices of $\beta$ in Table 8.
Ablation Study on Successive Elimination of Knowledge Set
The full results with Successive Elimination (SE) enabled or disabled are shown in Table 9.
C.2.3 Case Study
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF LaW ($\beta=0.2$, RI) 0.9158 0.2445 0.5082 10.86 0.5053 0.5059
LaW ($\beta=0.2$) 0.1258 0.1102 0.4708 13.04 0.4941 0.4831
zsRE LaW ($\beta=0.2$, RI) 0.8845 0.3274 0.5049 10.77 0.5081 0.5059
LaW ($\beta=0.2$) 0.0008 0.0008 0.4628 14.34 0.4924 0.4800
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF $\beta=1.05\beta_{0}$ 0.1266 0.1070 0.4743 12.49 0.5038 0.4950
$\beta=1.1\beta_{0}$ 0.1155 0.0995 0.4708 12.93 0.5017 0.4917
$\beta=1.2\beta_{0}$ 0.0965 0.0853 0.4553 14.10 0.4941 0.4829
$\beta=1.5\beta_{0}$ 0.0655 0.0587 0.4314 16.31 0.4819 0.4672
$\beta=0.1$ 0.4318 0.2401 0.5063 10.82 0.5055 0.5069
$\beta=0.2$ 0.1258 0.1102 0.4708 13.04 0.4941 0.4831
$\beta=0.5$ 0.0242 0.0220 0.3169 36.23 0.4209 0.4176
zsRE $\beta=1.05\beta_{0}$ 0.0074 0.0060 0.5127 10.74 0.5114 0.5096
$\beta=1.1\beta_{0}$ 0.0050 0.0039 0.5108 10.86 0.5079 0.5118
$\beta=1.2\beta_{0}$ 0.0008 0.0010 0.5073 11.06 0.5064 0.5126
$\beta=1.5\beta_{0}$ 0.0000 0.0003 0.4945 12.02 0.4960 0.5126
$\beta=0.1$ 0.0198 0.0166 0.5096 10.68 0.5097 0.5108
$\beta=0.2$ 0.0008 0.0008 0.4628 14.34 0.4924 0.4800
$\beta=0.5$ 0.0000 0.0000 0.2806 46.81 0.4242 0.4212
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF LaW 0.1091 0.0905 0.4741 12.73 0.5021 0.4909
w/o SE 0.1345 0.1137 0.4905 12.20 0.4843 0.5034
zsRE LaW 0.0058 0.0043 0.5114 10.81 0.5087 0.5114
w/o SE 0.0322 0.0254 0.5160 10.52 0.5103 0.5143
Wiki-Latest LaW 0.1926 0.1735 0.4657 13.27 0.4865 0.4975
w/o SE 0.2398 0.2038 0.4788 12.40 0.4940 0.5097

### v3

@@ 6.2.1 Small-Scale Knowledge Washing
We first test the performances of our method on the small-scale knowledge-washing tasks, i.e., forgetting the knowledge in zsRE and CounterFactual. We report the results in Table 1. As shown in the table, our method can achieve the best performance concerning the cleanness of knowledge washing (measured by Accuracy and QA-F1-Score) while maintaining performance levels comparable to the original model on reasoning tasks. As the dataset scale is not large, it is shown in the table that WOH and SeUL, two fine-tuning-based methods achieve some certain extent of knowledge washing and also successfully maintain the model’s original ability, although there is already sign of performance degradation as shown in the dataset CounterFactual (See the performances of SeUL on GPT2-XL). Meanwhile, the method FT-UL could not achieve reasonable results as the reverse training objective is overly fragile to the training without more complicated regularization. We also show the generated examples for visualization in Appendix C.3.6. To show the generalization of the knowledge unlearning abilities of our method, we conduct additional experiments using paraphrased queries from the zsRE and MCF datasets. The results are shown in Appendix 12.
@@ 6.2.2 Large Scale Knowledge Washing
Original 1.0000 0.3734 0.5105 1.0000 0.2553 0.6560
FT 0.0446 0.0256 0.3305 0.0159 0.0115 0.4867
MEMIT 0.2972 0.2342 0.5029 0.2536 0.0753 0.6436
ME-FT 0.0000 0.0000 0.1978 0.0000 0.0000 0.1716
FT-UL 0.0000 0.0000 0.1681 0.0000 0.0000 0.1669
WOH 0.4672 0.2227 0.2910 0.0009 0.0000 0.1728
SeUL 0.0000 0.0000 0.1647 0.0004 0.0000 0.1695
LaW 0.1926 0.1735 0.4832 0.1385 0.0846 0.6387
6.2.3 Unrelated Knowledge Preservation
W/zsRE W/CF W/Wiki
MEMIT 0.091 0.071 0.075
LaW 0.085 0.076 0.074
Table 3: The QA-F1 score of the model after washing some knowledge on 1000 examples extracted from Wikipedia. We evaluate the model after washing each dataset with MEMIT and LaW. The QA-F1 score of the base model GPT2-XL is 0.085. Here “W/” means “Washing”, “CF” and “Wiki” refer to “CounterFactual” and “Wiki-Latest”.
To evaluate the preservation of unrelated knowledge during the unlearning process, we create a new evaluation set containing 1,000 facts extracted from Wikipedia. The dataset is constructed using the same procedure as Wiki-Latest, as described in Section6.1. We focus on comparing the MEMIT method with our proposed approach, LaW (LaW), and present the results in Table 3. The results indicate that LaW performs comparably to MEMIT in retaining unrelated knowledge. Following Meng et al. (2023), for dataset zsRE and CounterFactual, we also include the analysis with the neighborhood prompts, which are prompts of similar forms as the target knowledge that needs to be unlearned, but contain different and unrelated knowledge. The analysis is shown in Appendix C.3.4.
6.2.4 Model Fluency Analysis
zsRE CounterFactual Wiki-Latest
GPT2-XL 2.66 2.66 2.66
MEMIT 2.67 2.70 2.72
SeUL 2.69 10.82 96.95
LaW 2.68 2.73 2.75
GPT-J-6B 2.31 2.31 2.31
MEMIT 2.31 2.33 2.37
SeUL 2.34 2.33 121.94
LaW 2.32 2.35 2.42
Table 4: Model Fluency Analysis.
To study whether the unlearning algorithms affect the model’s fluency, we sample 1000 examples from the snapshot CC-MAIN-2024-10 in dataset fineweb-edu (Penedo et al., 2024), and check the log perplexity of these models on the sampled subset. The results are reported in Table 4. From this table, we can observe that MEMIT and LaW barely affect the perplexity, whereas the unlearning baseline could drastically affect the perplexity, especially during large-scale settings (Wiki-Latest).
@@ 6.3 Ablation Study
We aim to explore the effects of the practical considerations described in Section 5.1. We put the experiments of Successive Elimination of Knowledge Set in Appendix C.3.5.
Ablation Study on Initialization of $\hat{\Delta}$. We compare the performance of LaW on the dataset zsRE and CounterFactual with model GPT2-XL between using random initialization and using the initialization from MEMIT. For random initialization, we sample a matrix matching the dimension of $W_{0}$ (in Eq.(2)), filled with independent Gaussian random variables scaled by a factor of $0.001$: $\Delta_{0}=0.001\cdot\mathcal{N}(0,I)$. The results are reported in Table 5 (full table in Appendix C.3.5). When initializing from Gaussian distribution, we do not have reference $\beta_{0}$ as in the initialization from MEMIT, so we choose the constant $\beta=0.2$. Similarly, we also set $\beta=0.2$ when using MEMIT initialization. The table shows the MEMIT initialization can boost the performance drastically. The reason might be the optimization easily achieves local minimum when using random optimization.
Ablation Study of Choices of $\beta$. As shown in Eq.(10), the hyper-parameter $\beta$ can control the tradeoff between washing the knowledge in $K_{w}$ and maintaining the original knowledge in $K$ (which may also be related to the model’s reasoning ability, as we find that when this term is large the model’s reasoning ability may degrade drastically). To study the effects of different $\beta$, we choose the setting of dataset zsRE and CounterFactual with the model GPT2-XL to study the effects of different $\beta$. The results are reported in Table 6 (full table in Appendix C.3.5). From the table, we can see that as $\beta$ increases, the knowledge is washed more thoroughly and the reasoning abilities are also dropping, showing the tradeoff between knowledge washing and maintaining reasoning abilities. We can also find that setting $\beta$ according to $\beta_{0}$ can achieve better performances (see the performance comparison between $\beta=1.2\beta_{0}$ and $\beta=0.2$ on the dataset CounterFactual), which demonstrates the necessity of setting different $\beta$ for different layers.
@@
GPT2-XL 1.0000 0.3734 0.5105 1.0000 0.3734 0.5105
$\beta=1.05\beta_{0}$ 0.0074 0.0060 0.5112 0.1266 0.1070 0.4910
$\beta=1.1\beta_{0}$ 0.0050 0.0039 0.5105 0.1091 0.0905 0.4881
$\beta=1.2\beta_{0}$ 0.0008 0.0010 0.5088 0.0965 0.0853 0.4774
$\beta=1.5\beta_{0}$ 0.0000 0.0003 0.5010 0.0655 0.0587 0.4602
$\beta=0.1$ 0.0198 0.0166 0.5100 0.4318 0.2401 0.5062
$\beta=0.2$ 0.0008 0.0008 0.4784 0.1258 0.1102 0.4827
$\beta=0.5$ 0.0000 0.0000 0.3753 0.0242 0.0220 0.3851
@@ 7 Conclusion, Limitation, and Future Work
Ethics Statement
Our research focuses on developing LaW, a method for large-scale knowledge washing in Large Language Models (LLMs), aiming to remove sensitive, private, or copyrighted information while preserving the models’ reasoning capabilities. We acknowledge the ethical considerations associated with both the presence of such information in LLMs and the processes involved in unlearning it.
Data Privacy and Compliance: The datasets used for unlearning in our experiments are derived from publicly available sources including zsRE (Levy et al., 2017), CounterFactual (Meng et al., 2022), and Wikipedia triplets, and do not contain personal or sensitive information about individuals.
Ethical Compliance: Throughout this study, we have adhered to the ICLR Code of Ethics. We conducted our research with integrity, respecting all applicable laws and ethical standards, and carefully considered the broader societal implications of our work.
Reproducibility Statement
We make sure the results are producible. We provide a clear experimental setup in Section  6.1. We provide our code as supplementary material to ensure the reproducibility.
@@ Appendix A Mathematical Details of Preliminary
As demonstrated by MEMIT (Meng et al., 2023), the objective is to adjust factual associations stored within the MLP layers of transformer-based, decoder-only large language models. The conditional distribution of the next token $x_{t}$, given by language model $G$, relies on the sequence of previous tokens:
@@
In MEMIT (Meng et al., 2023), $W_{out}^{l}$, denoted as $W_{0}$, can act as key-value memories, associating input keys $k_{i}\triangleq k_{i}^{l}$ with corresponding values $v_{i}\triangleq v_{i}^{l}$ (Geva et al., 2020). If $W_{out}^{l}$ is dimensionally defined as $d_{1}\times d_{2}$ and stores $n$ memories, with $u$ new edits, then to modify the MLP layer $W_{out}^{l}$ (i.e., the matrix $W_{0}$), the following delta matrix $\Delta$ is solved:
@@
This process enables the editing of an MLP layer within the transformer $G$ to incorporate new relational data, following the solution of the equation for each $K_{e}$ and $V_{e}$ from the editing requests. In the MEMIT approach (Meng et al., 2023), $KK^{T}$ is pre-estimated and represented as $\lambda C_{0}$, where $C_{0}$ is the average covariance matrix of $K$ and $\lambda$ is a hyper-parameter typically on the order of 10,000.
When performing extensive model editing, modifying only one layer may lead to robustness issues, while a more stable model can be achieved by minimizing the magnitudes of parameter changes (Zhu et al., 2020). Consequently, MEMIT proposes modifying multiple layers to distribute the editing impact more broadly (Meng et al., 2023). This method involves spreading the residual $R=V_{e}-W_{0}K_{e}$ across several layers. Let $L$ represent the index of the deepest layer requiring modification such that the output of this layer transitions from $W_{0}K_{e}$ to $V_{e}$. Define $\mathcal{R}$ as the set of layer indices $\{L-|\mathcal{R}|+1,\ldots,L\}$ that require edits. For each layer $l$ within $\mathcal{R}$, the necessary adjustments to the weights $W^{l}$ are given by:
@@ Appendix B Implementation Details
For the baselines, we train GPT-J-6B with LoRA (Hu et al., 2021). we put the configurations as below:
@@
    MEMIT. This method has a hyperparameter $\lambda$ when estimating $KK^{T}=\lambda C$ where $C$ is the average variable calculated on a large dataset (see the details in Meng et al. (2023)). The configurations of $\lambda$ in different settings are shown in Table 7. We found that with these configurations the model can achieve good knowledge-washing accuracy while mostly maintaining the model’s reasoning ability (minimal performance degradation on reasoning tasks.)
@@
    WOH. We first train the reinforced model on the sentences formed from the triplets $\mathcal{E}_{w}$ with the learning rate set as 1e-6 for 1 epoch, then we adopt the objective Eq.(1) from the paper Eldan & Russinovich (2023) to update the target model. During the second stage of training, we set the learning rate as 5e-5 and train the model for 1 epoch.
@@ Appendix C Additional Experiments
C.1 Choices of Models
Our current implementation is built on top of the MEMIT repository, which provides effective implementations only for GPT-2 and GPT-J. While we attempted to adapt MEMIT for LLaMA (specifically llama2-7b), the efficacy score did not exceed 0.75, compared to 0.96+ for GPT-2 and GPT-J (shown in Meng et al. (2023)). We assume this could result from configuration issues, such as selecting the appropriate layers to edit, tuning the value of of Eq.(15) of Meng et al. (2023), or estimating the covariance matrix ($\mathbb{E}_{k}[kk^{T}]$ ) using a more suitable dataset (it might need to be more aligned with LLaMA’s pretraining set rather than the Wikitext used in [1], for this part, please see Appendix B.3 in Meng et al. (2023)). Resolving these challenges would require significant adjustments unrelated to our proposed method. For these reasons, we chose to focus on GPT-2 and GPT-J in our experiments. We believe this decision allows us to present a fair and focused evaluation of LAW, while LLaMA integration remains a promising direction for future work.
C.2 Descriptions of the Reasoning Datasets
We conduct the reasoning experiments on three datasets: Lambda_openai (Radford et al., 2019; Paperno et al., 2016), HellaSwag (Zellers et al., 2019), and Arc_Easy (Clark et al., 2018). The descriptions of these three datasets are as follows:
@@ C.2 Descriptions of the Reasoning Datasets
    Lambda_openai (Paperno et al., 2016): The LAMBADA dataset tests computational text understanding via a word prediction task. It features narrative texts where models must use broad context to predict the final word, rather than just the last sentence. The dataset includes an original test split and translations in German, Spanish, French, and Italian.
@@
    HellaSwag (Zellers et al., 2019): The HellaSwag dataset is a benchmark designed for evaluating commonsense natural language inference (NLI) capabilities. It challenges models to complete sentences in a way that aligns with human common sense. The dataset prompts computational models to predict plausible sentence endings, testing their understanding of everyday scenarios and contexts.
@@
    ARC_Easy (Clark et al., 2018): The ARC_Easy dataset is a subset of the ARC dataset, featuring grade-school level multiple-choice science questions that are less challenging compared to the full set. It includes questions that were correctly answered by standard algorithms.
@@
Ours 0.1385 0.0846 0.6567 4.76 0.6452 0.5951
C.3 Additional Experimental Results
C.3.1 Overall Performance Comparison
The overall performance comparisons with the performances on two other reasoning benchmark on zsRE, CounterFactual and Wiki-Latest are shown in Table 8, Table 9 and Table 10, respectively.
C.3.2 Additional Experiments on GSM8k
In this section, we add the experiments on GSM-8K. We choose the strongest baseline in Table 1 and Table 2, which is MEMIT, and study the effects these knowledge-washing algorithms would have on the datasets GSM-8k. The results are shown in Table 11.
MCF flexible-extract strict-match
GPT2-XL 0.0205 $\pm$ 0.0039 0.0121 $\pm$ 0.0030
GPT2-XL-MEMIT 0.0182 $\pm$ 0.0037 0.0099 $\pm$ 0.0027
GPT2-XL-LaW 0.0197 $\pm$ 0.0038 0.0121 $\pm$ 0.0030
GPT-J-6B 0.0425 $\pm$ 0.0056 0.0349 $\pm$ 0.0051
GPT-J-6B-MEMIT 0.0440 $\pm$ 0.0056 0.0364 $\pm$ 0.0052
GPT-J-6B-LaW 0.0394 $\pm$ 0.0054 0.0334 $\pm$ 0.0049
zsRE flexible-extract strict-match
GPT2-XL 0.0205 $\pm$ 0.0039 0.0121 $\pm$ 0.0030
GPT2-XL-MEMIT 0.0190 $\pm$ 0.0038 0.0083 $\pm$ 0.0025
GPT2-XL-LaW 0.0205 $\pm$ 0.0039 0.0106 $\pm$ 0.0028
GPT-J-6B 0.0425 $\pm$ 0.0056 0.0349 $\pm$ 0.0051
GPT-J-6B-MEMIT 0.0394 $\pm$ 0.0054 0.0356 $\pm$ 0.0051
GPT-J-6B-LaW 0.0376 $\pm$ 0.0051 0.0336 $\pm$ 0.0049
Wiki-Latest flexible-extract strict-match
GPT2-XL 0.0205 $\pm$ 0.0039 0.0121 $\pm$ 0.0030
GPT2-XL-MEMIT 0.0167 $\pm$ 0.0035 0.0083 $\pm$ 0.0025
GPT2-XL-LaW 0.0212 $\pm$ 0.0040 0.0099 $\pm$ 0.0027
GPT-J-6B 0.0425 $\pm$ 0.0056 0.0349 $\pm$ 0.0051
GPT-J-6B-MEMIT 0.0440 $\pm$ 0.0056 0.0379 $\pm$ 0.0053
GPT-J-6B-LaW 0.0379 $\pm$ 0.0051 0.0326 $\pm$ 0.0047
Table 11: Performance (flexible-extract and strict-match) on the dataset GSM-8K after washing the knowledge in MCF, zsRE and Wiki-Latest.
C.3.3 Additional Experiments on Paraphrased Queries
To study the generalization of knowledge unlearning, we choose the strongest baseline in Table 1 and Table 2 and conduct the experiments using the paraphrased queries from zsRE and MCF datasets. The results are shown in Table 12.
Model zsRE Acc zsRE QA-F1 CounterFactual Acc CounterFactual QA-F1
GPT2-XL 0.6790 0.1371 0.4632 0.1308
MEMIT 0.0710 0.0506 0.1958 0.1349
LAW 0.0140 0.0104 0.1274 0.1027
GPT-J-6B 0.6966 0.1537 0.5059 0.1507
MEMIT 0.0231 0.0102 0.2623 0.1037
LAW 0.0036 0.0026 0.1492 0.0668
Table 12: Performance on the paraphrased set of zsRE and MCF datasets.
C.3.4 Analysis of Retaining Unrelated Knowledge
To address the concern about retaining unrelated knowledge, we conducted additional experiments using neighborhood prompts from the zsRE and CounterFactual datasets. These prompts probe the model for unrelated knowledge after washing specific facts. The results are summarized in Table 13.
zsRE CounterFactual
GPT2-XL 0.0150 0.0626
GPT2-XL-MEMIT 0.0157 0.0697
GPT2-XL-LaW 0.0139 0.0734
GPT-J-6B 0.0270 0.0671
GPT-J-6B-MEMIT 0.0308 0.0648
GPT-J-6B-LaW 0.0254 0.0579
Table 13: QA-F1-Score on neighborhood prompts
C.3.5 Ablation Study
Ablation Study on Initialization of $\hat{\Delta}$
Ablation Study of Choices of $\beta$
We put the full results of different choices of $\beta$ in Table 15.
Ablation Study on Successive Elimination of Knowledge Set
In our practical considerations, before modifying every layer, we find the facts in the knowledge set that the model can answer correctly and perform the knowledge washing on the selected knowledge set. To study the effects of this technique (denoted as SE), we conduct experiments with and without SE and report the results in Table 16 The results show that the algorithm can achieve a much cleaner washing with SE enabled, at the expense of slightly affecting the reasoning abilities.
C.3.6 Case Study
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF LaW ($\beta=0.2$, RI) 0.9158 0.2445 0.5082 10.86 0.5053 0.5059
LaW ($\beta=0.2$) 0.1258 0.1102 0.4708 13.04 0.4941 0.4831
zsRE LaW ($\beta=0.2$, RI) 0.8845 0.3274 0.5049 10.77 0.5081 0.5059
LaW ($\beta=0.2$) 0.0008 0.0008 0.4628 14.34 0.4924 0.4800
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF $\beta=1.05\beta_{0}$ 0.1266 0.1070 0.4743 12.49 0.5038 0.4950
$\beta=1.1\beta_{0}$ 0.1155 0.0995 0.4708 12.93 0.5017 0.4917
$\beta=1.2\beta_{0}$ 0.0965 0.0853 0.4553 14.10 0.4941 0.4829
$\beta=1.5\beta_{0}$ 0.0655 0.0587 0.4314 16.31 0.4819 0.4672
$\beta=0.1$ 0.4318 0.2401 0.5063 10.82 0.5055 0.5069
$\beta=0.2$ 0.1258 0.1102 0.4708 13.04 0.4941 0.4831
$\beta=0.5$ 0.0242 0.0220 0.3169 36.23 0.4209 0.4176
zsRE $\beta=1.05\beta_{0}$ 0.0074 0.0060 0.5127 10.74 0.5114 0.5096
$\beta=1.1\beta_{0}$ 0.0050 0.0039 0.5108 10.86 0.5079 0.5118
$\beta=1.2\beta_{0}$ 0.0008 0.0010 0.5073 11.06 0.5064 0.5126
$\beta=1.5\beta_{0}$ 0.0000 0.0003 0.4945 12.02 0.4960 0.5126
$\beta=0.1$ 0.0198 0.0166 0.5096 10.68 0.5097 0.5108
$\beta=0.2$ 0.0008 0.0008 0.4628 14.34 0.4924 0.4800
$\beta=0.5$ 0.0000 0.0000 0.2806 46.81 0.4242 0.4212
Lambda_openai hellaswag arc_easy
[…]
GPT2-XL 1.0000 0.3734 0.5121 10.63 0.5089 0.5105
CF LaW 0.1091 0.0905 0.4741 12.73 0.5021 0.4909
w/o SE 0.1345 0.1137 0.4905 12.20 0.4843 0.5034
zsRE LaW 0.0058 0.0043 0.5114 10.81 0.5087 0.5114
w/o SE 0.0322 0.0254 0.5160 10.52 0.5103 0.5143
Wiki-Latest LaW 0.1926 0.1735 0.4657 13.27 0.4865 0.4975
w/o SE 0.2398 0.2038 0.4788 12.40 0.4940 0.5097